# Machine Unlearning: A Comprehensive Survey

### v2

@@ 3.2. Machine Unlearning Evaluation Metrics
Kullback–Leibler divergence (KLD). KLD is commonly used to measure the divergence between two probability distributions, often assessing the distance between retrained and unlearned models. In Bayes-based or Markov chain Monte Carlo-based unlearning methods (Nguyen et al., 2020), researchers utilize KLD (Joyce, 2011) to optimize approximate models, employing it to measure the distance between two probability distributions. Recent unlearning studies have also used KLD to estimate the unlearning effect by comparing the distributions of retrained and unlearned models (Nguyen et al., 2020).
Evaluation Metric based on Privacy Leakage. Since membership inference attacks (Chen et al., 2021) can decide whether a sample was utilized for training a model, recently, some works have leveraged this property to verify if unlearning mechanisms remove the specific data. Some studies (Hu et al., 2022) even proposed to backdoor the unlearning samples for initial model training and then attack the unlearned model. If the unlearned model is still backdoored, this proves that the unlearning algorithm cannot unlearn samples effectively. Conversely, if the backdoor trigger cannot attack the unlearned model, it proves that the unlearning algorithm is effective. Similar methods were also used in (Thudi et al., 2022a; Ma et al., 2022) to evaluate the unlearning effect.
3.3. Majority of Tools Used in Unlearning
@@
4. Centralized Machine Unlearning
In this section, we classify existing centralized unlearning methods by their inherent mechanism and designed purposes and present the corresponding detailed techniques.
@@ 4.1. Unlearning Solution Categories
From the former introduction, we know that naive retraining is the most effective manner to realize machine unlearning. However, it is inefficient because it requires storing the entire original dataset and retraining the model from scratch, which consumes significant storage and computational resources, especially in deep learning scenarios. Therefore, researchers tried to design effective and efficient unlearning mechanisms. Exact unlearning was proposed to reduce the computation cost by splitting training sub-models based on pre-divided data sub-sets (Bourtoule et al., 2021). It also divides the stochasticity and the incrementality into sub-models. It can unlearn an exact model by ensembling the consisting submodels, but it still needs to store all the split data subsets for fast retraining. Another approach is approximate unlearning (Sekhari et al., 2021). It can reduce both storage and computation costs because it unlearns only based on the erased data. However, it is difficult for approximate unlearning methods to control the accuracy degradation due to the challenges in estimating stochasticity and incrementality during the training process. Most of them bounded their estimation to avoid dramatic accuracy reduction.
@@ 4.1.1. Exact Unlearning
Exact unlearning could also be called fast retraining, whose basic idea is derived from naive retraining from scratch. Following the background introduction of unlearning, we know the learning and unlearning algorithm, $\mathcal{A}(D)$ and $\mathcal{U}$, based on the trainset $D$ and erased dataset $D_{e}\subseteq D$, respectively. If $\mathcal{U}(\cdot)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\setminus D_{e})\in\mathcal{H}$ and $\mathcal{U}(M,D,D_{e})\in\mathcal{H}$ is absolutely guaranteed. However, naive retraining involves high computation and storage costs, especially for deep learning models and complex datasets (Thudi et al., 2022a). Unlike naive retraining, which relies on the whole remaining dataset, exact unlearning tries to retrain a sub-model only using a subset of the remaining dataset to reduce calculation cost. A general operation of exact unlearning is that they first divide the dataset into several small sub-sets. Then, they transform the learning process by ensembling the sub-models trained with each sub-set as the final model (Bourtoule et al., 2021; Cao and Yang, 2015). So that when an unlearning request comes, they are just required to retrain the sub-model corresponding to the sub-set containing the erased data. They then ensemble the retrained sub-model and other sub-models as the unlearned model.
As we introduced above, exact unlearning aims to mitigate the computation cost when retraining a new model by transforming the original learning algorithms into an ensembling form. It divides the stochasticity and incrementality into several sub-models to reduce their influence. However, to some extent, they sacrificed the storage cost because they needed to store the whole training dataset in a divided form. In (Cao and Yang, 2015), Cao and Yang transformed the traditional ML algorithms into a summation form. They are only required to update several summations when an unlearning requirement comes, ensuring the method runs faster than retraining from scratch. SISA (Bourtoule et al., 2021) is a representative exact unlearning algorithm, which splits the full training dataset into shards and trained models separately in each shard. For unlearning, they simply need to retrain the shard that includes the erased data. Study (Wu et al., 2022a) proposed a framework that precisely models the impact of individual training sample on the model concerning various performance criteria and removes the impact of samples that are required to be removed. Golatkar et al. (Golatkar et al., 2021) proposed an unlearning method on deep networks, splitting the trained model into two parts. The core part based on the data will not be deleted, and the unlearning part with the erased data will be unlearned with parameters bound. These methods are efficient in computation, but they sacrifice the storage space to store the intermediate training parameters of different slices and the related training sub-sets.
@@ 4.1.2. Approximate Unlearning
Many methods were proposed to implement approximate unlearning efficiently and effectively. The popular solutions are certified-removal (Guo et al., 2020) and Bayes-based mechanisms (Nguyen et al., 2020), which are introduced in technical detail in Section 4.2. Here, we briefly summarize current studies about approximate unlearning. A certified-removal mechanism (Guo et al., 2020) was an approximate unlearning method similar to differential privacy. Guo et al. (Guo et al., 2020) defined the $\epsilon-$approximate unlearning, which makes sure that the model after and before unlearning must be $\epsilon$-indistinguishable as the definition in DP. The difference between $\epsilon-$approximate unlearning and $\epsilon$-DP is that the mechanism $\mathcal{A}$ on differential privacy is needed never to memorize the data in the first place, which is impossible in machine unlearning. The machine learning model does not learn anything from the training dataset if $\mathcal{A}$ is differentially private (Bourtoule et al., 2021). A similar solution to (Guo et al., 2020) is introduced in (Sekhari et al., 2021), which erases the influence of the specified samples from the gradients during unlearning. To mitigate accuracy degradation, model parameter updates are constrained by a differential unlearning definition.
Different from indistinguishable unlearning, other methods used the Bayes theorem to unlearn an approximate posterior. In (Nguyen et al., 2020), Nguyen et al. employed Bayesian inference to unlearn an approximate model using the erased data. Moreover, Fu et al. (Fu et al., 2022) developed a similar unlearning method, which uses Markov chain Monte Carlo (MCMC) methods. In (Nguyen et al., 2022a), the authors also explained the effectiveness of the approximate unlearning method from the perspective of MCMC. Nevertheless, since those techniques are approximately unlearning the contribution of all the erasing data, including the inputs and labels, they inevitably decrease the model accuracy to some extent after unlearning.
Main Challenges of Approximate Unlearning. As we discussed at the beginning, there are three main challenges in machine unlearning. In centralized scenarios, these studies mainly aim at solving the basic machine unlearning problem, which unavoidably faces three challenges: stochasticity of training, incrementality of training, and catastrophe of unlearning. The exact unlearning methods extend the retraining idea, which avoids facing these challenges but consumes lots of storage costs. The approximate unlearning methods face these challenges directly, and we here list the relevant work about how to estimate the contribution of erased samples to overcome the stochasticity and incrementality of training, and how to prevent unlearning catastrophe.
  • To overcome the stochasticity and incrementality challenges when estimating the unlearning influence, one popular strategy is based on the first-order and second-order influence function (Basu et al., 2020), which is calculated based on the perturbation theory (Avrachenkov et al., 2013).
@@
4.2. Detailed Techniques of Centralized Unlearning
Table 2. Popular Unlearning Techniques
@@ 4.2. Detailed Techniques of Centralized Unlearning
SUMMIT (Zhang et al., 2024) Approximate unlearning Graph nodes Certified Data Removal Multi-Objective Optimization 2024
@@
This section presents the popular techniques used in existing unlearning methods, in both exact and approximate unlearning studies. Exact unlearning extends the idea of retraining and tries to reduce the computation cost of unlearning. Approximate unlearning was proposed to find a way to reduce computation and storage consumption together. The dominant studies are summarized in Table 2, where the primary technique used in exact unlearning is split learning. Two primary techniques used in approximate unlearning are certified data removal and Bayesian-based unlearning. The introduction of these techniques will primarily follow the sequence of the unlearning techniques presented in Fig. 1.
@@ 4.2.2. Certified Data Removal
The aforementioned methods are trying to address the basic unlearning problem from a certified data removal perspective. Neel et al. (Neel et al., 2021) extended the definition of unlearning to include updates, which encompass both ”add” ($D\cup\{z\}$) and ”delete” ($D\backslash\{z\}$), where $z=(x,y)$ and $z\in\mathcal{Z}$ is a data point. They follow the definition in (Guo et al., 2020) and define similar $(\epsilon,\delta)$-indistinguishability. Furthermore, they extend $(\epsilon,\delta)$-indistinguishability to both ”add” and ”delete” updates, which can also be denoted as $(\epsilon,\delta)$-publishing.
@@ 4.2.3. Bayesian-based Unlearning
4.2.4. Graph Unlearning
We introduce graph unlearning as a representative kind of irregular data unlearning. In (Chien et al., 2022a; Chen et al., 2022b; Cong and Mahdavi, 2022; Zhang et al., 2024), researchers extend regular data machine unlearning to a graph data scenario. Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig. 7. Therefore, Chien et al. (Chien et al., 2022a) proposed node unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC). Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees. They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
5. Federated Unlearning
We have introduced popular centralized machine unlearning in the former section. Their main purpose is to reduce the calculation and storage costs and the accuracy degradation caused by unlearning. Recently, we noticed that some researchers paid attention to realizing unlearning in other scenarios, such as in federated learning (FL) scenarios. In this section, we will introduce the literature about federated unlearning.
FL was initially introduced to protect the privacy of participating clients during the machine learning training process in distributed settings. All participants will only upload their locally trained model parameters instead of their sensitive local data to the FL server during model training processes. Therefore, in a federated learning scenario, limited access to the dataset will become a unique challenge when implementing unlearning.
Since the local data cannot be uploaded to the federated learning (FL) server side, most federated unlearning methods try to erase a certain client’s contribution from the trained model by storing and estimating the contribution of uploaded parameters. In this situation, they can implement federated unlearning without interacting with the client, shown as the server-side federated unlearning in Fig. 8 (a). The two representative methods are (Liu et al., 2021; Wu et al., 2022b). Liu et al. (Liu et al., 2021) proposed “FedEraser” to sanitize the impact of a FL client on the global FL model. In particular, during FL training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates. Based on these operations, they reconstructed the unlearned FL model instead of retraining a new model from scratch. However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data. It is unsuitable for a client who wants to unlearn a small piece of his data. Study (Wu et al., 2022b) tried to erase a client’s influence from the FL model by removing the historical updates from the global model. They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
Wang et al. (Wang et al., 2022a) explored the problem of how to selectively unlearn a category from a trained CNN classification model in FL. The biggest challenge is that the local data utilized for the FL global model training cannot be accessed globally. Therefore, they explored the inner influence of each channel and observed that various channels have distinct impacts on different categories. They proposed a method that does not require accessing the data used for training and retraining from scratch to cleanly scrub the information of particular categories from the global FL model. A method called TF-IDF was introduced to quantize the class discrimination of channels. Then, they pruned the channels with high TF-IDF scores on the erased classifications to unlearn them. Evaluations of CIFAR10 and CIFAR100 demonstrate the unlearning effect of their method.
Different from unlearning a whole client’s influence and unlearning a class, Liu et al. (Liu et al., 2022a; Wang et al., 2023c) investigated how to unlearn data samples in FL, shown as the client-side federated unlearning in Fig. 8 (b). In (Liu et al., 2022a), they first defined a federated unlearning problem and proposed a fast retraining method to withdraw the influence of data from the FL model. Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods and the first-order Taylor approximate method (Liu et al., 2022a). They utilized the practical Fisher Information Matrix to model the Hessian matrix at a low cost. When unlearning, they subtract the estimation matrix from gradients with a bound. In (Wang et al., 2023c), they implement federated unlearning based on Bayesian inference, and they propose a parameters self-sharing method to reduce the model utility degradation.
@@ 6. Unlearning Evaluation and Verification
L2-norm The parameters differences between the retrained and unlearned models $||\theta_{1}-\theta_{2}||$ To evaluate the indistinguishability between two models (Wu et al., 2020)
@@
Membership inference Recall (##\##detected objects /#absent#/\#/ #erased objects) To verify if the erased sample is unlearned by the model (Graves et al., 2021)
@@
EMA Ensembled Membership Auditing: Ensemble multiple membership metrics and utilizes Kolmogorov-Smirno (KS) statistical tools to obtain a final auditing score To verify if a unlearned model mem -orizes a query dataset (Liu and Tsaftaris, 2020)
@@
6.1. The Unlearning Verification and Auditing Methods
@@
Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effect is also an important problem in machine unlearning. In Section 3, we have introduced the L2-norm (Wu et al., 2020), KLD (Nguyen et al., 2020), and privacy leakage as the unlearning verification metrics. The common metrics also include accuracy in assessing the performance of the unlearned model and running time in evaluating the unlearning efficiency. We have listed recent evaluation and verification metrics in Table 3, where the popular experimental datasets are furthermore presented. Since accuracy, running time, L2-norm, and KLD are fixed evaluation metrics, we do not discuss them in detail here. In this section, we introduce some new evaluation metrics that recent studies tailored to unlearning. Most of them are based on privacy leakage.
In (Liu and Tsaftaris, 2020), Liu and Tsaftaris indicated that the data auditing problem in machine unlearning is becoming popular. They used a calibration dataset and the Kolmogorov-Smirno (KS) (Demidenko, 2004) distance to confirm if an erased dataset is removed. Their challenge was that the method would fail when the calibration dataset was out of high quality, or the query samples were similar to the training dataset. The EMA (Huang et al., 2021a) method was proposed to overcome these limitations when auditing data erasure from a trained model. EMA is a 2-step data removal auditing method that ensembles several metrics and statistical tools. In the first step of EMA, the auditor uses various metrics to infer if the model remembers the samples of the query dataset. EMA aggregates a final auditing score in the second step by ensembling many membership metrics and statistical tools. They verified whether a trained model forgets query data by the final auditing score. They conducted the experiments using standard datasets, including MNIST, SVHN, and the Chest X-ray dataset.
@@
Inspired by backdoor attacks in ML, Hu et al. (Hu et al., 2022) proposed Membership Inference via Backdooring (MIB). MIB leverages the property of backdoor attacks that backdoor triggers will misadvise the trained model to predict the backdoored sample to other wrong classes. The main idea of MIB is that the user proactively adds the trigger to her data when publishing them online so that she can implement the backdoor attacks to determine if the model has been trained using her dataset. MIB evaluates the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model. Sequentially, they observed the results between the targeted model and clean models to infer whether the model was trained based on the backdoored samples.
@@
A similar membership-inference-based method was proposed in (Ma et al., 2022). Ma et al. (Ma et al., 2022) verified the effect of unlearning methods by the proposed forgetting rate (FR) metric. The evaluation metrics were defined using the observation of membership inference. Suppose $D_{e}$ is the erased dataset; the FR of an unlearning method is denoted as $FR=\frac{AF-BF}{BT}$. In their FR definition, $BF$ and $AF$ are the samples in $D_{e}$ that are predicted as false by a membership inference attack before and after machine unlearning operations. $BT$ is the sample size in $D_{e}$, which is indicated to be correct by a membership inference attack before machine unlearning. According to the definition, $FR$ presents an instinctive evaluation of how many data points are altered from member to non-member by unlearning. If an unlearning method achieves that $AF>BF$ on the condition that $BT>0$, it means this method is effective. By contrast, the unlearning will be meaningless. They also proposed an unlearning method called Forsaken and evaluated Forsaken’s performance on eight standard benchmark datasets. Forsaken’s main idea is to generate masked gradients continuously, apply them to the model’s neurons, and model them to forget the contribution of the specific samples. To implement this process, they introduced a mask gradient generator. The proposed mask gradient generator could adjust the volume to prevent unlearning catastrophes by observing the updated model.
@@
Sommer et al. (Sommer et al., 2022) introduced “Athena”, which leverages the property of backdooring techniques to verify the effectiveness of unlearning. Athena effectively and confidently certifies whether the data is deleted from an unlearning method. Thus, it provides a basis for quantitatively inferring unlearning. In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is unlearned.
@@
Data Type Name of Datasets Feature Dimension #. Smaples Task Type Employed by
@@
YELP2018 $5$ $1,561,406$ Recommendation (Chen et al., 2022a)
Movielens-1m $4$ $1,000,209$ Recommendation (Chen et al., 2022a)
Movielens-10m $4$ $10,000,054$ Recommendation (Chen et al., 2022a)
Text AG News (Zhang et al., 2015) $3$ $127,600$ Text classification (Wu et al., 2020)
RCV1 (Glewis et al., 2004) $3$ $804,414$ Text classification (Wu et al., 2020)

### v3

@@ 3.2. Machine Unlearning Evaluation Metrics
Kullback–Leibler divergence (KLD). KLD is commonly used to measure the divergence between two probability distributions, often assessing the distance between retrained and unlearned models. In Bayes-based or Markov chain Monte Carlo-based unlearning methods (Nguyen et al., 2020), researchers utilize KLD (Joyce, 2011) to optimize approximate models, employing it to measure the distance between two probability distributions. Recent unlearning studies have also used KLD to estimate the unlearning effectiveness by comparing the distributions of retrained and unlearned models (Nguyen et al., 2020).
Evaluation Metric based on Privacy Leakage. Since membership inference attacks (Chen et al., 2021) can decide whether a sample was utilized for training a model, recently, some works have leveraged this property to verify if unlearning mechanisms remove the specific data. Some studies (Hu et al., 2022) even proposed to backdoor the unlearning samples for initial model training and then attack the unlearned model. If the unlearned model is still backdoored, this proves that the unlearning algorithm cannot unlearn samples effectively. Conversely, if the backdoor trigger cannot attack the unlearned model, it proves that the unlearning algorithm is effective. Similar methods were also used in (Thudi et al., 2022a; Ma et al., 2022) to evaluate the unlearning effectiveness.
3.3. Tools Used in Unlearning
@@
4. Traditional Centralized Machine Unlearning
In this section, we classify existing traditional centralized unlearning methods by their inherent mechanism and designed purposes and present the corresponding detailed techniques.
@@ 4.1. Unlearning Solution Categories
From the former introduction, we know that naive retraining is the most effective manner to realize machine unlearning. However, it is inefficient because it requires storing the entire original dataset and retraining the model from scratch, which consumes significant storage and computational resources, especially in deep learning scenarios. Therefore, researchers tried to design effective and efficient unlearning mechanisms, and two representative solutions are exact unlearning and approximate unlearning.
@@ 4.1.1. Exact Unlearning
Exact unlearning is also called fast retraining, whose basic idea is derived from naive retraining from scratch. Following the background of unlearning, we know the learning and unlearning algorithm, $\mathcal{A}(D)$ and $\mathcal{U}$, based on the trainset $D$ and erased dataset $D_{e}\subseteq D$, respectively. If $\mathcal{U}(\cdot)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\setminus D_{e})\in\mathcal{H}$ and $\mathcal{U}(M,D,D_{e})\in\mathcal{H}$ is absolutely guaranteed. However, naive retraining involves high computation and storage costs, especially for deep learning models and complex datasets (Thudi et al., 2022a). Unlike naive retraining, which relies on the whole remaining dataset, exact unlearning tries to retrain a sub-model only using a subset of the remaining dataset to reduce calculation cost. A general operation of exact unlearning is that they first divide the dataset into several small sub-sets. Then, they transform the learning process by ensembling the sub-models trained with each sub-set as the final model (Bourtoule et al., 2021; Cao and Yang, 2015). So that when an unlearning request comes, they are just required to retrain the sub-model corresponding to the sub-set containing the erased data. They then ensemble the retrained sub-model and other sub-models as the unlearned model.
Exact unlearning aims to mitigate the computation cost when retraining a new model by transforming the original learning algorithms into an ensembling form. It divides the stochasticity and incrementality into several sub-models to reduce their influence. However, to some extent, they sacrificed the storage cost because they needed to store the whole training dataset in a divided form.
  • In (Cao and Yang, 2015), Cao and Yang transformed the traditional ML algorithms into a summation form. They are only required to update several summations when an unlearning requirement comes, ensuring the method runs faster than retraining from scratch.
  • SISA (Bourtoule et al., 2021) is a representative exact unlearning algorithm, which splits the full training dataset into shards and trained models separately in each shard. For unlearning, they simply need to retrain the shard that includes the erased data.
  • Study (Wu et al., 2022a) proposed a framework that precisely models the impact of individual training sample on the model concerning various performance criteria and removes the impact of samples that are required to be removed.
  • Golatkar et al. (Golatkar et al., 2021) proposed an unlearning method on deep networks, splitting the trained model into two parts. The core part based on the data will not be deleted, and the unlearning part with the erased data will be unlearned with parameters bound.
These methods are efficient in computation, but they sacrifice the storage space to store the intermediate training parameters of different slices and the related training sub-sets.
@@ 4.1.2. Approximate Unlearning
Many methods were proposed to implement approximate unlearning efficiently and effectively. The popular solutions are certified-removal (Guo et al., 2020) and Bayes-based mechanisms (Nguyen et al., 2020), which are introduced in technical detail in Section 4.2. Although those techniques are approximately unlearning the contribution of all the erasing data, including the inputs and labels, they inevitably decrease the model accuracy to some extent after unlearning.
Main Challenges of Approximate Unlearning. In centralized scenarios, researchers aiming at solving the basic machine unlearning problem will unavoidably face three challenges: stochasticity of training, incrementality of training, and catastrophe of unlearning. The exact unlearning methods extend the retraining idea, which avoids facing these challenges but consumes lots of storage costs. The approximate unlearning methods face these challenges directly, and we here list the relevant work about how to estimate the contribution of erased samples to overcome the stochasticity and incrementality of training, and how to prevent unlearning catastrophe.
  • To overcome the stochasticity and incrementality challenges when estimating the unlearning influence, one popular strategy is based on the first-order and second-order influence function (Basu et al., 2020), which is calculated based on the perturbation theory (Avrachenkov et al., 2013). At the same time, since classical influence functions are best justified under smooth and strongly convex objectives, whereas deep neural networks usually involve highly non-convex loss landscapes, the local Taylor approximation and inverse-Hessian estimation can become unstable (Koh and Liang, 2017; Basu et al., 2020).
@@
4.2. Detailed Techniques of Traditional Centralized Unlearning
Table 3. Traditional Unlearning Techniques
@@ 4.2. Detailed Techniques of Traditional Centralized Unlearning
SUMMIT (Zhang et al., 2024a) Approximate unlearning Graph nodes Certified Data Removal Multi-Objective Optimization 2024
@@
This section presents the popular techniques used in existing unlearning methods, in both exact and approximate unlearning studies. Exact unlearning extends the idea of retraining and tries to reduce the computation cost of unlearning. Approximate unlearning was proposed to find a way to reduce computation and storage consumption together. The dominant studies are summarized in Table 3, where the primary technique used in exact unlearning is split learning. Two primary techniques used in approximate unlearning are certified data removal and Bayesian-based unlearning.
@@ 4.2.2. Certified Data Removal
[…]
@@ 5. Unlearning Evaluation and Verification
Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effectiveness is also an important problem in machine unlearning. In this section, we will introduce some common and basic unlearning verification methods and evaluation datasets.
@@
L2-norm The parameters differences between the retrained and unlearned models $\theta_{1}-\theta_{2}||$ To evaluate the indistinguishability between two models (Wu et al., 2020)
@@
Membership inference Recall (#\#detected objects /#/\#erased objects) To verify if the erased sample is unlearned by the model (Graves et al., 2021)
@@
EMA Ensembled Membership Auditing: Ensemble multiple membership metrics and utilizes Kolmogorov-Smirnov (KS) statistical tools to obtain a final auditing score To verify if an unlearned model memorizes a query dataset (Liu and Tsaftaris, 2020)
@@
5.1. The Unlearning Verification Methods
@@
In Section 3, we have introduced the L2-norm (Wu et al., 2020), KLD (Nguyen et al., 2020), and privacy leakage as the unlearning verification metrics. The common metrics also include accuracy in assessing the performance of the unlearned model and running time in evaluating the unlearning efficiency. We have listed these evaluation metrics in Table 4. Here, we will introduce some new evaluation metrics that recent studies tailored to unlearning.
@@
Attack-based verification. Inspired by backdoor attacks in ML, Hu et al. (Hu et al., 2022) proposed Membership Inference via Backdooring (MIB). MIB leverages the property of backdoor attacks that backdoor triggers will misadvise the trained model to predict the backdoored sample to other wrong classes. The main idea of MIB is that the user proactively adds the trigger to her data when publishing them online so that she can implement the backdoor attacks to determine if the model has been trained using her dataset. MIB evaluates the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model. Although MIB is effective to verify the unlearning of backdoored samples, it is hard to directly verify the unlearning of benign samples.
@@
A similar membership-inference-based method was proposed in (Ma et al., 2022). Ma et al. (Ma et al., 2022) verified the effectiveness of unlearning methods by the proposed forgetting rate (FR) metric. The evaluation metrics were defined using the observation of membership inference. Suppose $D_{e}$ is the erased dataset; the FR of an unlearning method is denoted as $FR=\frac{AF-BF}{BT}$. In their FR definition, $BF$ and $AF$ are the samples in $D_{e}$ that are predicted as false by a membership inference attack before and after machine unlearning operations. $BT$ is the sample size in $D_{e}$, which is indicated to be correct by a membership inference attack before machine unlearning. According to the definition, $FR$ presents an instinctive evaluation of how many data points are altered from member to non-member by unlearning. If an unlearning method achieves that $AF>BF$ on the condition that $BT>0$, it means this method is effective. By contrast, the unlearning will be meaningless. The membership inference is an important verification tool, but failing an MIA is a necessary but not sufficient condition for successful unlearning.
@@
Sommer et al. (Sommer et al., 2022) introduced “Athena”, which leverages the property of backdooring techniques to verify the effectiveness of unlearning. Athena effectively and confidently certifies whether the data is deleted from an unlearning method. Thus, it provides a basis for quantitatively inferring unlearning. In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is unlearned. Like the MIB method, Athena is also hard to directly verify the unlearning of benign data, and the backdooring technique should be processed before the original model training.
Model-centric (influence or model-difference) audits. Another line of research views the difference between the model before and after unlearning as a meaningful source of evidence. EMU trains reconstruction models over simulated model differences to infer attributes of the deleted data, thereby measuring residual privacy leakage without depending on explicit backdoor design (Wang et al., 2025b). Building on this perspective, TAPE focuses on tailored posterior differences and enhances auditing with data perturbation and influence-based partitioning, enabling a more fine-grained assessment of how much private information about forgotten samples can still be recovered across different unlearning settings (Wang et al., 2025c). Similarly, TruVRF adopts a non-invasive, model-centric perspective by leveraging model sensitivity to verify unlearning at class-, volume-, and sample-level granularities, extending model-change-based auditing toward finer-grained white-box verification without relying on explicit backdoor design (Zhou et al., 2025). These approaches move auditing beyond direct output probing and toward more principled evidence grounded in influence patterns and model-change signatures.
@@
Data Type Name of Datasets Feature Dimension #. Samples Task Type Employed by
@@
YELP2018 55 $1,561,406$ Recommendation (Chen et al., 2022a)
Movielens-1m 44 $1,000,209$ Recommendation (Chen et al., 2022a)
Movielens-10m 44 $10,000,054$ Recommendation (Chen et al., 2022a)
Text AG News (Zhang et al., 2015) 33 $127,600$ Text classification (Wu et al., 2020)
RCV1 (Glewis et al., 2004) 33 $804,414$ Text classification (Wu et al., 2020)